# HANS-Net: Hyperbolic Convolution and Adaptive Temporal Attention for Accurate and Generalizable Liver and Tumor Segmentation in CT Imaging

### v1

@@ 5.3   Comparison With Existing Literature
Previous models such as MSML-AttUNet [11] attained performances of 87.74% on combined datasets like LiTS and 3DIRCADb. Although models like EG-UNETR [8] exhibited commendable performance at 85.01% and S2DA-Net [7] demonstrated moderate results with 80.0%, HANS-Net presents a marked improvement. Models such as MCDA [12] and G-Unet [9] recorded lower scores, suggesting constraints in generalization or segmentation precision. Even methodologies employing private datasets, such as MCFMFNet [5], did not exceed the 90% threshold. The comparative analysis highlights the strength and practical effectiveness of HANS-Net, showing that it has a distinct advantage over the existing SOTA models. Table 4 provides a detailed comparison between the proposed approach and contemporary state-of-the-art techniques.
@@
[8] EG-UNETR LiTS 84.45%
[4] SBCNet LiTS 82.50%
[9] G-Unet LiTS 72.90%
@@ 5.4   Cross-Dataset Evaluation
To thoroughly evaluate the generalizability of our proposed HANS-Net model, we conducted cross-dataset evaluations (see Table 5) on the 3D-IRCADb-01 dataset, assessing segmentation performance for liver and tumor in different acquisition protocols and anatomical variations.
Table 5: Comparative performance of HANS-Net and state-of-the-art segmentation models on the 3D-IRCADb-01 dataset for liver and tumor segmentation. The best and second-best results for each dataset are bolded and underlined, respectively.
[…]
U-Net [29] Liver 91.25 84.72 1.48 15.28
Tumor 75.84 62.93 2.85 37.07
Mean 83.55 73.83 2.17 26.18
U-Net++ [30] Liver 92.47 86.18 1.35 13.82
Tumor 77.62 65.24 2.61 34.76
Mean 85.05 75.71 1.98 24.29
Attention U-Net [16] Liver 93.18 87.53 1.19 12.47
Tumor 78.95 67.46 2.43 32.54
Mean 86.07 77.50 1.81 22.51
TransUNet [32] Liver 93.94 88.67 1.08 11.33
Tumor 80.16 69.85 2.27 30.15
Mean 87.05 79.26 1.68 20.74
H-DenseUNet [31] Liver 92.83 86.91 1.24 13.09
Tumor 78.31 66.82 2.52 30.18
Mean 85.57 76.87 1.88 21.64
HANS-Net (ours) Liver 94.82 90.85 0.97 9.15
Tumor 80.08 69.74 2.08 30.26
Mean 87.45 80.30 1.53 19.71
Figure 8: Comparative performance of our proposed HANS-Net model and baseline methods for liver and tumor segmentation on the 3D-IRCADb-01 dataset.
Table 6: Quantitative uncertainty and correlation metrics.
@@
Pearson Correlation 0.999 0.995 0.997
Spearman Correlation 0.998 0.993 0.995
@@
Table 7: Ablation study on the LiTS dataset with averaged liver-tumor metrics. The configurations include components such as Hyperbolic Convolution (HC), Wavelet Decomposition Module (WDM), Adaptive Temporal Attention (ATA), Synaptic Plasticity Module (SPM), Implicit Neural Representation (INR), and Uncertainty Estimator (UE).
@@
HC WDM ATA SPM INR UE Train Val Train Val Train Val Train Val Epc×$\times$×T (M)
A1 87.55 84.84 79.45 76.20 1.45 1.75 20.55 23.80 100$\times$230s 0.89M
A2 89.60 87.23 82.58 79.18 1.24 1.46 17.43 20.83 100$\times$235s 0.98M
A3 90.79 88.36 84.35 81.13 1.05 1.21 15.65 18.88 100$\times$238s 1.05M
A4 91.89 89.63 86.03 82.78 0.94 1.07 13.98 17.23 100$\times$240s 1.12M
A5 92.60 90.38 87.33 84.00 0.83 0.97 12.68 16.00 100$\times$242s 1.15M
A6 93.18 91.03 88.20 84.85 0.77 0.91 11.80 15.15 100$\times$243s 1.17M
A7 93.86 91.47 89.17 85.86 0.66 0.89 10.83 14.77 100$\times$245s 1.17M
The proposed model achieves strong generalization performance on the 3D-IRCADb-01 dataset, with a Dice score of 94.82%, IoU of 90.85%, ASSD of 0.97 mm, and VOE of 9.15% for liver segmentation. For tumor segmentation, HANS-Net attains 80.08% Dice, 69.74% IoU, 2.08 mm ASSD, and a VOE of 30.26%. Averaging across organs, the model achieves 87.45% Dice, 80.30% IoU, 1.53 mm ASSD, and 19.71% VOE. Compared to state-of-the-art methods, HANS-Net consistently outperforms all baselines. U-Net [29] achieves 83.55% mean Dice and 73.83% IoU, while U-Net++ [30] and Attention U-Net [16] improve to 85.05% and 86.07% Dice, respectively. H-DenseUNet [31] records 85.57% Dice and 76.87% IoU, whereas TransUNet [32] achieves 87.05% Dice and 79.26% IoU. HANS-Net provides a +0.40% Dice and +1.04% IoU improvement over TransUNet, the next best model. It also achieves the lowest ASSD (1.53 mm) and VOE (19.71%) among all methods. Although H-DenseUNet reports the lowest tumor VOE of 30.18%, our HANS-Net model demonstrates superior balance across all key metrics. Figure 8 visualizes the performance of comparative segmentation compared to current methods.
@@ 5.5   Uncertainty-Aware Performance Evaluation
As shown in Figure 9, the uncertainty heatmaps clearly highlight regions of lower model confidence, particularly at organ boundaries and areas of tumor heterogeneity. Quantitative evaluation in all test samples provides consistent uncertainty measures with mean values of 0.619 for liver and tumor predictions; and for both predictions, the uncertainty range was observed between 0.595 and 0.637 (see Table 6). The narrow range of 0.042 indicates stable model behavior across different cases. These values represent the uncertainty characteristics observed in our liver and tumor segmentation task, where the model successfully distinguishes between confident and uncertain predictions. While our model demonstrates effective uncertainty quantification for liver and tumor segmentation, the determination of acceptable uncertainty thresholds is application-specific and would require domain expert validation for different medical imaging tasks and clinical contexts.
@@ 5.6   Volumetric Correlation Analysis
As shown in Table 6, the HANS-Net model demonstrates robust clinical validity through its volumetric measurements, with near-perfect Pearson’s correlation coefficients of 0.999 for liver and 0.995 for tumor volumes alongside Spearman’s correlations of 0.998 and 0.993, respectively. All correlations show statistical significance with p-values below $1\times 10^{-100}$, while maintaining precise volumetric measurements with mean absolute errors of 25.8 mL for liver and 13.6 mL for tumors, corresponding to relative differences of just 2.0% and 6.6%. The model’s mean predicted volumes of 1362.6 mL for liver and 252.3 mL for tumors closely approximate the ground truth values of 1359.9 mL and 252.0 mL, respectively. Figure 10 demonstrates the model’s measurement reliability through the close alignment of data points with the identity line.
@@ 5.7   Ablation Study
We conducted a systematic ablation study to analyze the effect of progressive enhancement achieved by HANS-Net. In this section, we explore the main components, including hyperbolic convolution (HC), wavelet decomposition module (WDM), adaptive temporal attention (ATA), synaptic plasticity module (SPM), implicit neural representation (INR) and uncertainty estimator (UE). The configuration details are reported in Table 7.
The baseline encoder-decoder architecture (A1) establishes the foundation with conventional convolutions and shows a notable train-validation gap that motivates our architectural innovations. Then, instead of traditional Euclidean convolutions, we employ hyperbolic convolutions (A2) to capture hierarchical anatomical structures in curved space, delivering a 2.39% Dice improvement and 16.57% ASSD reduction while enhancing IoU by approximately 3%.
Next, rather than relying on single-scale processing, we introduce the wavelet decomposition module (A3), which enables multifrequency analysis for fine-detail detection. This addition provides an additional 1.13% dice gain and an improvement of 17. 12% ASSD with minimal computational overhead. Furthermore, in place of standard feature aggregation, the adaptive temporal attention module (A4) is incorporated to model slicewise correlations, achieving a 1.27% Dice enhancement and an 11.57% reduction in ASSD, thus improving boundary precision.
Furthermore, unlike static feature learning, we integrate a synaptic plasticity module (A5), introducing bioinspired adaptation mechanisms. This module contributes an additional 0.75% Dice improvement and 9.35% ASSD reduction while helping minimize train-validation discrepancies. Furthermore, instead of using discrete feature representations, we employ implicit neural representations (A6), which enable continuous coordinate-based learning. This technique delivers a 0.65% dice gain and a 6.19% ASSD improvement, all while maintaining exceptional parameter efficiency.
@@ 6   Discussion
Our proposed model comprises several advanced modules that enhance its robustness and overall performance. The incorporation of hyperbolic convolutions for hierarchical geometric representation, wavelet-based decomposition for multiscale segmentation, synaptic plasticity for adaptive feature enhancement, and an implicit neural representation branch for precise boundary modeling collectively contributed to enhanced segmentation accuracy. To further quantify the confidence of prediction, an uncertainty-aware Monte Carlo dropout is used. We further integrated adaptive temporal attention to enhance the consistency of inter-slice without decreasing efficiency.
@@
The integration of implicit neural representation contributed to the generation of smooth and precise organ boundaries, which significantly improved the quality of segmentation. Moreover, our approach effectively captures both global structures and fine local details, which played a crucial role in boosting the Dice scores. To further improve performance, we incorporated a lightweight self-attention mechanism with reduced channel dimensions, utilized a multilayer perceptron for learning adaptive temporal weighting, and optimized the query-key-value attention architecture to enhance consistency across slices. The addition of such other advanced modules, such as hyperbolic convolution, uncertainty estimator, and implicit neural representation, greatly increased the robustness of the model, leading to superior segmentation performance.
@@
The proposed HANS-Net model demonstrates significant segmentation performance in the LiTS data set for both liver and tumor structures. With respect to liver segmentation, the model achieved a Dice score of 96.67% and an IoU of 93.79%, which signifies an excellent score for segmenting the liver. The model also showed a low ASSD of 0.54 mm and a minimal VOE of 6.21%, indicating precise boundary definition and low volumetric error. Although the results for tumor segmentation are slightly lower due to increased structural complexity, they remain robust, with a Dice score of 89.84%, IoU of 82.39%, ASSD of 0.89 mm and VOE of 17.61%. In summary, the model achieved a mean Dice coefficient of 93.26%, IoU of 88.09%, ASSD of 0.72 mm and VOE of 11. 91%, thus confirming its efficiency in segmenting large and small anatomical structures.
@@
The robustness of HANS-Net is further evidenced by its stable convergence seen in both the training and validation curves, which shows no major signs of overfitting. Interestingly, liver Dice scores rapidly surpassed 90% within the first 20 epochs, while tumor Dice scores increased more slowly, reaching about 80% in epoch 80. This trend highlights a hierarchical learning process that first captures broad anatomical structures before focusing on tumor boundary details. Metrics such as ASSD and VOE showed consistent decreases, pointing to precise segmentation and a growing confidence in the model. The evaluation of the cross-dataset on the 3D-IRCADb-01 dataset confirms the impressive generalizability of the model, with a mean Dice coefficient of 87.45%, and an IoU of 80. 30%, and low ASSD (1.53 mm) and VOE (19.71%).
@@
Compared to recent SOTA models, HANS-Net achieves superior segmentation performance, attaining a mean Dice score of 93.26%, whereas most existing methods struggle to surpass the 90% threshold. For instance, MSML-AttUNet and FUF-TransUnet reached Dice scores of 87.74% and 89.77%, respectively, while models such as MCDA and G-UNet performed even lower. Among pre-trained architectures, only TransUNet exceeded 90%, yet it remained approximately 3% lower than HANS-Net, highlighting the advantage of our framework in both accuracy and structural consistency.
@@
As shown in Table 7, an ablation study reveals the progressive benefit of our proposed modules. The base configuration starts with a Dice score of 87.55%, which rises to 89.60% after integrating hyperbolic convolution. Incorporating all components—including synaptic plasticity, temporal attention, and uncertainty modeling—culminates in a final Dice score of 93.86%, demonstrating the additive effectiveness of each module.
While HANS-Net demonstrates promising results, it has several limitations. First, in cross-dataset evaluation on the 3D-IRCADb-01 dataset, the model’s performance showed a modest decline, with the Dice score dropping to 80.08% and the IoU to 69.74%. This decrease is likely due to the considerable variability in tumor appearance, comprising differences in intensity, texture, size, and anatomical location as well as the irregular and often indistinct boundaries of tumors. Second, although the implicit representation branch enhances boundary continuity, it introduces additional computational overhead and increases inference time. Third, the current study focuses exclusively on liver and tumor segmentation from CT scans. The generalization of model to other imaging modalities such as MRI or PET has not yet been assessed. Finally, the model has been evaluated only on a single-organ scenario. Its adaptability to multi-organ segmentation tasks remains unexplored, which is essential for real-world clinical deployment where multi-organ analysis is often required.

### v2

@@ 5.3 Comparison With Existing Literature
Previous models, such as MSML-AttUNet [16], attained performances of 87.74% on combined datasets like LiTS and 3DIRCADb. Although models like EG-UNETR [12] exhibited commendable performance at 85.01% and S2DA-Net [11] demonstrated moderate results with 80.0%, HANS-Net presents a marked improvement. Models such as MCDA [17] and G-Unet [13] recorded lower scores, suggesting constraints in generalization or segmentation precision. Even methodologies employing private datasets, such as MCFMFNet [8], did not exceed the 90% threshold. The comparative analysis highlights the strength and practical effectiveness of HANS-Net, showing that it has a distinct advantage over the existing SOTA models. Table 3 provides a detailed comparison between the proposed approach and the SOTA techniques.
@@
[23] DPC-Net LiTS 76.40%
@@ 5.4 Cross-Dataset Evaluation
Table 4: Segmentation performance of HANS-Net on the AMOS 2022 dataset across 15 abdominal organs. The abbreviated organs are: Right Kidney (R. Kid.); Left Kidney (L. Kid.); Stomach (Stom.); Pancreas (Panc.); Right Adrenal Gland (R. Adr.); Left Adrenal Gland (L. Adr.); Duodenum (Duod.); Urinary Bladder (Blad.); Prostate (Prost.).
Metric Spleen R. Kid. L. Kid. Gallbladder Esophagus Liver Stom. Aorta Postcava Panc. R. Adr. L. Adr. Duod. Blad. Prost. Mean
Dice (%) 81.37 83.36 80.85 88.56 81.51 94.17 85.52 80.96 94.13 86.23 87.65 87.90 80.72 83.53 79.89 85.09
IoU (%) 75.93 78.84 73.16 75.58 75.06 80.12 74.42 69.83 88.94 79.48 84.32 75.24 71.05 77.90 70.08 76.66
ASSD (mm) 20.63 26.31 20.37 15.96 24.94 8.35 16.72 25.89 9.65 18.88 17.75 13.53 21.39 19.44 32.49 19.49
VOE (%) 24.07 21.16 26.84 24.42 24.94 19.88 25.58 30.17 11.06 20.52 15.68 24.76 28.95 22.10 29.92 23.34
To further assess the model’s cross-dataset generalization capabilities, we evaluated the proposed HANS-Net model on the AMOS 2022 dataset, which contains 15 annotated abdominal organs.
Figure 8: Cross-dataset prediction visualizations from our HANS-Net model on AMOS 2022 dataset.
Figure 8 and Table 4 present the overall segmentation performance of our proposed model on the AMOS 2022 test set. The results demonstrate robust performance across all anatomical structures, with our model achieving an overall average Dice score of 85.09%, IoU of 76.66%, ASSD of 19.49 mm, and VOE of 23.34%. Notably, the model performed exceptionally well on large solid organs such as the liver (94.17% Dice) and the postcava (94.13% Dice), as well as the gallbladder (88.56% Dice). Even for challenging structures like the duodenum and prostate/uterus, which are typically difficult to segment due to their complex anatomical variations and low contrast, our model achieved Dice scores of 80.72% and 79.89%, respectively. The low ASSD values for liver (8.35 mm) and postcava (9.65 mm) further validate the effectiveness of our implicit neural representation in maintaining precise boundary delineation.
Table 5: Statistical consistency and agreement metrics between HANS-Net predicted and ground-truth liver and tumor volumes, including correlation coefficients, bias, mean absolute errors, normalized errors, 95% limits of agreement, and uncertainty ranges.
@@
Pearson Correlation ($r$) 0.999 0.995 0.997
Spearman Correlation ($\rho$) 0.998 0.993 0.995
Coefficient of Determination ($R^{2}$) 0.998 0.990 0.994
@@
Normalized MAE (% of GT) 1.9 5.4 3.65
95% Limits of Agreement (mL) [–47.87, +53.27] [–26.36, +26.9] N/A
Uncertainty Range 0.595 – 0.637 0.619
Table 6: Ablation study on the LiTS dataset with averaged liver-tumor metrics. The configurations include components such as Hyperbolic Convolution (HC), Wavelet Decomposition Module (WDM), Adaptive Temporal Attention (ATA), Synaptic Plasticity Module (SPM), Implicit Neural Representation (INR), and Uncertainty Estimator (UE). The configurations are experimented with for 100 epochs, and the inference speed for the configurations is provided in seconds in the Inf. Time column.
@@
[…]
@@ 5.5 Uncertainty-Aware Performance Evaluation
As shown in Figure 9, the uncertainty heatmaps highlight regions of lower model confidence, particularly at organ boundaries and areas of tumor heterogeneity. Quantitative evaluation in all test samples provides consistent uncertainty measures with mean values of 0.619 for liver and tumor predictions; and for both predictions, the uncertainty range was observed between 0.595 and 0.637 (see Table 5). The narrow range of 0.042 indicates stable model behavior across different cases. These values represent the uncertainty characteristics observed in our liver and tumor segmentation task, where the model successfully distinguishes between confident and uncertain predictions. While our model demonstrates effective uncertainty quantification for liver and tumor segmentation, the determination of acceptable uncertainty thresholds is application-specific and would require domain expert validation for different medical imaging tasks and clinical contexts.
@@ 5.6 Volumetric Correlation Analysis
As shown in Table 5, the HANS-Net model demonstrates robust clinical validity through its volumetric measurements, with near-perfect Pearson’s correlation coefficients of 0.999 for liver and 0.995 for tumor volumes alongside Spearman’s correlations of 0.998 and 0.993, respectively. All correlations show statistical significance with p-values below $1\times 10^{-100}$.
These strong correlations are supported by the Coefficients of Determination, $R^{2}$ (0.998 for liver and 0.990 for tumor) and show that over 99% of the variance in the ground-truth volumes is captured by the model predictions. The mean predicted volumes of 1362.6 mL for liver and 252.3 mL for tumors closely match the ground-truth means of 1359.9 mL and 252.0 mL, and this results in negligible Mean Bias Errors of +2.7 mL and +0.3 mL. The Mean Absolute Errors are 25.8 mL for liver and 13.6 mL for tumor, and they correspond to Normalized Errors of 1.9% and 5.4%. In addition to these findings, the reported 95% limits of agreement are [–47.9, +53.3] mL for liver and [–26.4, +27.0] mL for tumor, and most predictions remain within a narrow and interpretable range of the true volumes. These results collectively demonstrate that HANS-Net provides statistically reliable and clinically interpretable volumetric estimations. Figure 10 visualizes the model’s measurement reliability through the close alignment of data points with the identity line.
@@ 5.7 Ablation Study
We conducted a systematic ablation study to analyze the effect of progressive enhancement achieved by HANS-Net. In this section, we explore the main components, including hyperbolic convolution (HC), wavelet decomposition module (WDM), adaptive temporal attention (ATA), synaptic plasticity module (SPM), implicit neural representation (INR), and uncertainty estimator (UE). The configuration details are reported in Table 6.
Initially, we experimented with a standard 3D U-Net architecture (A0) that operates within Euclidean space, where feature correlations are defined through linear kernel operations and assumed flat spatial geometry. We observed that while the configuration was effective for local intensity modeling, this formulation could not accurately represent the curved anatomical surfaces commonly present in volumetric data. As a result, the Euclidean baseline exhibited limited geometric expressiveness in validation across the defined metrics despite having higher training performance. These outcomes indicate that Euclidean convolutions primarily learn low-level texture dependencies (which requires more time) rather than topological or structural consistency across slices. Thus, in the next configuration, our designed baseline encoder-decoder architecture (A1) adopts a hierarchical architecture consisting of four encoding and decoding stages. Each stage includes two $\times3$ convolutional layers followed by batch normalization and ReLU activation, starting with 32 channels and doubling at each subsequent level (32, 64, 128, 256). Downsampling is achieved using $\times2$ max-pooling, while upsampling utilizes transposed convolutions with concatenated skip connections. The output layer applies a $\times1$ convolution with sigmoid activation to generate pixel-wise predictions. Training was performed using the Adam optimizer (learning rate = 1e-4) with a combined Dice and Cross-Entropy loss and ReduceLROnPlateau scheduler for 100 epochs and a batch size of 16. This baseline establishes the foundation with conventional convolutions and shows a notable train-validation gap that motivates our architectural innovations. Then, instead of traditional Euclidean convolutions, we employ hyperbolic convolutions (A2) to capture hierarchical anatomical structures in curved space, delivering a 2.39% Dice improvement and 16.57% ASSD reduction while enhancing IoU by approximately 3%.
Next, rather than relying on single-scale processing, we introduce the wavelet decomposition module (A3), which enables multifrequency analysis for fine-detail detection. This addition provides an additional 1.13% dice gain and an improvement of 17. 12% ASSD with minimal computational overhead. Furthermore, in place of standard feature aggregation, the adaptive temporal attention module (A4) is incorporated to model slice-wise correlations, achieving a 1.27% Dice enhancement and an 11.57% reduction in ASSD, thus improving boundary precision.
Furthermore, unlike static feature learning, we integrate a synaptic plasticity module (A5), introducing bio-inspired adaptation mechanisms. This module contributes an additional 0.75% Dice improvement and 9.35% ASSD reduction while helping minimize train-validation discrepancies. Furthermore, instead of using discrete feature representations, we employ implicit neural representations (A6), which enable continuous coordinate-based learning. This technique delivers a 0.65% dice gain and a 6.19% ASSD improvement, all while maintaining exceptional parameter efficiency.
@@
To verify that performance gains are not artifacts of integration order, we conducted permutation experiments by altering the module sequence, such as incorporating WDM before HC or introducing INR earlier. Across five randomized permutations, the average validation Dice varied by only 0.31%, with ASSD fluctuations under 0.08 mm, indicating order-independent contributions. This stability arises because each component targets orthogonal aspects: HC refines geometric encoding, WDM captures multiscale textures, ATA enforces temporal coherence, SPM adapts features dynamically, INR provides continuous refinement, and UE quantifies confidence. Their modular design ensures minimal interference, confirming that improvements reflect genuine architectural benefits rather than sequence-dependent interactions.
Analyzing the computational parameter and inference time, we notice a gradual increase across the configurations that reflects the cumulative computational impact of each added module, corresponding to deeper feature hierarchies and denser inter-module communication. The introduction of hyperbolic convolutions and the wavelet decomposition module slightly extends forward-pass operations due to additional manifold projections and frequency decompositions, yet these remain linear in parameter growth. Subsequent modules, such as adaptive temporal attention and synaptic plasticity, add minor tensor–matrix interactions that marginally increase runtime but significantly enhance representational adaptability. The near-constant epoch increment from 230s to 245s indicates that the additional parameters (0.28M overall) primarily improve weight expressiveness rather than computational depth, meaning the model achieves richer spatiotemporal encoding without a proportional rise in latency.
@@ 6 Discussion
Our proposed model comprises several advanced modules that enhance its robustness and overall performance. The incorporation of hyperbolic convolutions [31] for hierarchical geometric representation, wavelet-based [30] decomposition for multiscale segmentation, synaptic plasticity [32] for adaptive feature enhancement, and an implicit neural representation [19] branch for precise boundary modeling collectively contributed to enhanced segmentation accuracy. To further quantify the confidence of prediction, an uncertainty-aware Monte Carlo dropout is used [34]. We further integrated adaptive temporal attention to enhance the consistency of inter-slice without decreasing efficiency [20]. While our method operates on 2D slices for computational efficiency, volumetric continuity is preserved through multi-view extraction and adaptive temporal attention that explicitly models sequential relationships between adjacent slices. Since the liver channel encompasses both parenchyma and tumor, all liver-containing slices are evaluated regardless of tumor presence, maintaining sensitivity to potential false positives in tumor-free regions.
@@
The integration of implicit neural representation contributed to the generation of smooth and precise organ boundaries, which significantly improved the quality of segmentation [19]. Moreover, our approach effectively captures both global structures and fine local details, which played a crucial role in boosting the Dice scores. To further improve performance, we incorporated a lightweight self-attention mechanism with reduced channel dimensions, utilized a multilayer perceptron for learning adaptive temporal weighting, and optimized the query-key-value attention architecture to enhance consistency across slices [20]. The addition of such other advanced modules, such as hyperbolic convolution, uncertainty estimator, and implicit neural representation, greatly increased the robustness of the model, leading to superior segmentation performance.
@@
The proposed HANS-Net model demonstrates significant segmentation performance in the LiTS dataset [35] for both liver and tumor structures. With respect to liver segmentation, the model achieved a Dice score of 96.67% and an IoU of 93.79%, which signifies an excellent score for segmenting the liver. The model also showed a low ASSD of 0.54 mm and a minimal VOE of 6.21%, indicating precise boundary definition and low volumetric error. Although the results for tumor segmentation are slightly lower due to increased structural complexity, they remain robust, with a Dice score of 89.84%, IoU of 82.39%, ASSD of 0.89 mm, and VOE of 17.61%. In summary, the model achieved a mean Dice score of 93.26%, IoU of 88.09%, ASSD of 0.72 mm, and VOE of 11. 91%, thus confirming its efficiency in segmenting large and small anatomical structures.
@@
The robustness of HANS-Net is further evidenced by its stable convergence seen in both the training and validation curves, which show no major signs of overfitting. Interestingly, liver Dice scores rapidly surpassed 90% within the first 20 epochs, while tumor Dice scores increased more slowly, reaching about 80% in epoch 80. This trend highlights a hierarchical learning process that first captures broad anatomical structures before focusing on tumor boundary details. Metrics such as ASSD and VOE showed consistent decreases, pointing to precise segmentation and a growing confidence in the model. The evaluation of the cross-dataset on the AMOS 2022 [36] dataset confirms the impressive generalizability of the model, with a mean Dice score of 85.09%, an IoU of 76.66%, and low ASSD (19.49 mm) and VOE (23.34%).
@@
Compared to recent SOTA models, HANS-Net achieves superior segmentation performance, attaining a mean Dice score of 93.26%, whereas most existing methods struggle to surpass the 90% threshold. For example, MSML-AttUNet [16] reached Dice scores of 87.74%, respectively, while models such as MCDA [17] and G-UNet [13] performed even lower. Among pre-trained architectures, only TransUNet [2] exceeded 90%, yet it remained approximately 3% lower than HANS-Net, highlighting the advantage of our framework in both accuracy and structural consistency.
@@
As shown in Table 6, an ablation study reveals the progressive benefit of our proposed modules. The base configuration starts with a Dice score of 87.55%, which rises to 89.60% after integrating hyperbolic convolution. Incorporating all components, including synaptic plasticity, temporal attention, and uncertainty modeling, culminates in a final Dice score of 93.86%, demonstrating the additive effectiveness of each module. The inference time and model size analysis further confirm its suitability for clinical deployment. With only 0.28M additional parameters and a minimal increase of +15s in processing time, the model maintains practical efficiency and can be deployed in standard hospital GPU workstations without requiring specialized hardware.
While HANS-Net demonstrates promising results, it has several limitations. First, although our evaluation computes standard segmentation metrics (Dice, IoU, ASSD, VOE) at the slice level, future work should report these metrics aggregated at the per-volume level to align with the original LiTS benchmark protocol. Additionally, while processing resized 2D slices, the model utilizes dynamic sequential weighting to maintain volumetric continuity, a mechanism that could be further adapted to achieve complete 3D reconstruction. Second, in cross-dataset evaluation on the AMOS 2022 dataset, the model’s performance showed a moderate decline, with the Dice score dropping to 85.09% and the IoU to 76.66%. This decrease can be attributed to the high anatomical diversity of abdominal organs, variations in imaging modalities, and inter-organ boundary ambiguities present in the dataset. Third, although the implicit representation branch enhances boundary continuity, it introduces additional computational overhead and increases inference time.